In how many ways 6 boys can sit on a sofa, if only 3 seets are available?
Answer by rothauserc(4718)

6C3 = 6! / (3! * (6-3)!) = (6*5*4) / (3*2*1) = 5*4 = 20
:
**************************************
There are 20 ways 6 boys can be seated
**************************************
: